OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} (-1)^{n+1} \frac{n^{2} 2^{n}}{n!}\] ratio test?

OpenStudy (anonymous):

\[\frac{(n+1)^{2}2^{(n+1)}}{(n+1)!} * \frac{n!}{n^{2}2^{n}}\]

OpenStudy (turingtest):

\[(-1)^{n+2}{(n+1)^22^{n+1}\over(n+1)!}\cdot\frac1{(-1)^{n+1}}{n!\over n^22^n}\]

OpenStudy (anonymous):

it becomes \[\frac{n^2}{2n+2}\]

OpenStudy (anonymous):

TuringTest...where did you get the \[(-1)^{n+2}\] from?

OpenStudy (anonymous):

sorry flit up

OpenStudy (anonymous):

\[\frac{2n+2}{n^2}\]

OpenStudy (turingtest):

\[(-1)^{n+2}{(n+1)^22^{n+1}\over(n+1)!}\cdot\frac1{(-1)^{n+1}}{n!\over n^22^n}\]\[(-1){2n^2+4n+2\over n^2(n+1)}\]@MathSofiya the term after (-1)^(n+1) is (-1)^(n+2)

OpenStudy (anonymous):

oops!.... thanks

OpenStudy (turingtest):

\[(-1)^{n+2}{(n+1)^22^{n+1}\over(n+1)!}\cdot\frac1{(-1)^{n+1}}{n!\over n^22^n}\]\[(-1){2n^2+4n+2\over n^2(n+1)}=-2{\left({n+1}\over n^2\right)}\]does anybody disagree with that? cinar where did the - go ?

OpenStudy (turingtest):

whatever, it's absolute value anyway

OpenStudy (turingtest):

and it tends to zero

OpenStudy (anonymous):

yeah that's good..

OpenStudy (experimentx):

I think for alternating series ,,, proving the elements tends to zero as n tends to infinity and each term is decreasing term {of absolute sequence } would be enough. however absolute convergence implies convergence.

OpenStudy (experimentx):

check this out http://tutorial.math.lamar.edu/Classes/CalcII/AlternatingSeries.aspx

OpenStudy (anonymous):

why does my solution manual ignores the \[(-1)^{n+1}\] ?

OpenStudy (anonymous):

Turning Test please fan me you are AWESOME!!!!

OpenStudy (experimentx):

Hmm ... it's an alternating series, proving the convergence of absolute term ... also implies convergence

OpenStudy (anonymous):

exactly...that's what Paul online notes does too

OpenStudy (experimentx):

If \[ \frac{n^{2} 2^{n}}{n!}\] is convergent then \[ (-1)^n \frac{n^{2} 2^{n}}{n!} \] is obviously convergent. Though it can trouble you sometime . for eg \( \frac 1 n \) is divergent while \( (-1)^n \frac 1n \) is convergent.

OpenStudy (anonymous):

\[ b_{n}\] is \[\frac{n^{2} 2^{n}}{n!}\] so we can ignore \[(-1)^{n+1}\] when calculating convergence using the ratio test ...correct?

OpenStudy (experimentx):

Yes you can, but do not always use this method. Why? see my above post. it's better if you show \[ \frac{(n+1)^22^{n+1}}{(n+1)!} \leq \frac{(n)^22^{n}}{(n)! }\] and \[ \lim_{x \rightarrow \infty } \frac{(n)^22^{n}}{(n)! } = 0\]

OpenStudy (experimentx):

you cannot use that method on \[ \sum_{1}^\infty \frac{(-1)^n} n \]

OpenStudy (anonymous):

Ok \[\frac{(n+)^{2}2^{(n+1)}}{n^{2}2^{n}} * \frac{n!}{(n+1)!}\] I'm not the best at doing limits...so lets see the fraction on the right becomes \[\infty\]?

OpenStudy (anonymous):

plz don't give up on me just yet

OpenStudy (turingtest):

was in not you whom I talked to yesterday about dividing factorials?

OpenStudy (anonymous):

yes and the term become n+1

OpenStudy (anonymous):

or n in this case

OpenStudy (turingtest):

right, keep going

OpenStudy (anonymous):

ok n+1

OpenStudy (turingtest):

1/(n+1) in this case

OpenStudy (anonymous):

n-> \[\infty\] so 1/1

OpenStudy (turingtest):

no limits yet, we're not done simplifying

OpenStudy (experimentx):

\[ \frac{(n+1)^{2}2^{(n+1)}}{n^{2}2^{n}} * \frac{n!}{(n+1)!} = \left ( 1+ \frac 1n \right )^{2}2^{1} \frac{1}{n+1} \] taking limit you get zero which is less than 1 so the series converges absolutely.

OpenStudy (anonymous):

ok so I have \[ \frac{(n+)^{2}2^{(n+1)}}{n^{2}2^{n}}\] * \[\frac{1}{n+1}\]

OpenStudy (turingtest):

right, so how can you simplify this part\[\frac{(n+1)^{2}2^{(n+1)}}{n^{2}2^{n}}\]?

OpenStudy (experimentx):

Interesting value http://www.wolframalpha.com/input/?i=sum [n^2+2^n%2Fn!%2C+1%2C+infinity]

OpenStudy (turingtest):

^weird...

OpenStudy (experimentx):

this one is interesting too http://www.wolframalpha.com/input/?i=sum [%28-1%29^%28n%2B1%29+n^2+2^n%2Fn!%2C+1%2C+infinity]

OpenStudy (anonymous):

\[\frac{(n+1)^{2}2^{(n+1)}}{n^{2}2^{n}}\] = ok I can't do the elimination we did with factorials...so I don't really know.... Please give me hint

OpenStudy (turingtest):

never forget that\[\large\frac{x^a}{x^b}=x^{a-b}\]

OpenStudy (experimentx):

the fastest way to type latex is copying. Right click on latex ... show Math as Tex commands

OpenStudy (turingtest):

right click->show math as->tex commands then copy and paste, and enclose in brackets \[\.] (without the dot . )

OpenStudy (anonymous):

\[\frac{(n+1)^{2}2^{(n+1)}}{n^{2}2^{n}}\] = \[\frac{(n+1)^{2} 2}{n^{2}}\]

OpenStudy (turingtest):

yes :) times...?

OpenStudy (anonymous):

\[\frac{(n+1)^{2} 2}{n^{2}}\] * \[\frac{1}{n+1}\]

OpenStudy (anonymous):

i seeeeeeeee

OpenStudy (anonymous):

\[\frac{(n+1)2}{n^{2}}\]

OpenStudy (turingtest):

and the limit of that as n tends to infinity is...?

OpenStudy (anonymous):

1/ infty =0

OpenStudy (anonymous):

WOW!!! 43 MIN LATERRR!!!

OpenStudy (anonymous):

guess better later than never ey

OpenStudy (turingtest):

notice how the division makes the n drop out mostly in cases like this, so when you get the hang of it you'll see certain parts coming. for instance, that 2^(x+1)/2^x=2 is really normal in these problems, so next time I bet you'll do it a lot faster like how you learned factorials :)

OpenStudy (anonymous):

Thanks

OpenStudy (turingtest):

welcome!